# BATCLIP: Bimodal Online Test-Time Adaptation for CLIP

### v2

@@ Abstract
Although open-vocabulary classification models like Contrastive Language Image Pretraining (CLIP) have demonstrated strong zero-shot learning capabilities, their robustness to common image corruptions remains poorly understood. Through extensive experiments, we show that zero-shot CLIP lacks robustness to common image corruptions during test-time, necessitating the adaptation of CLIP to unlabeled corrupted images using test-time adaptation (TTA). However, we found that existing TTA methods have severe limitations in adapting CLIP due to their unimodal nature. To address these limitations, we propose BATCLIP, a bimodal online TTA method designed to improve CLIP’s robustness to common image corruptions. The key insight of our approach is not only to adapt the visual encoders for improving image features but also to strengthen the alignment between image and text features by promoting a stronger association between the image class prototype, computed using pseudo-labels, and the corresponding text feature. We evaluate our approach on benchmark image corruption datasets and achieve state-of-the-art results in online TTA for CLIP. Furthermore, we evaluate our proposed TTA approach on various domain generalization datasets to demonstrate its generalization capabilities.
@@ 1 Introduction
To address these core limitations, we propose, BATCLIP, a bimodal adaptation approach of CLIP for online TTA specifically, where both the visual and text encoders are jointly adapted by exploiting CLIP’s shared feature space of images and text. The overall objective is to achieve a strong alignment between image features and text features to enable more effective multimodal learning and adaptation. The adaptation procedure is two-fold: 1) Inspired by [54, 58] for efficient fine-tuning, we only adapt the norm layers, i.e., LayerNorm parameters of both encoders. However, such a model update does not consider the alignment of the encoder features. Therefore, to improve the alignment between class-specific visual and text features, we introduce a projection matching loss that maximizes the projection of the visual class prototypes with their corresponding text features. 2) To learn more discriminative visual features, we enhance the cosine distance between class prototypes, computed using pseudo-labels, to promote a more distinct separation in the image feature space. Our method is general-purpose, with no reliance on multiple prompt templates [36, 15, 34] or their ensembles, unlike VTE. We leverage batches of test samples for TTA, rather than focusing on single-image test adaptation, which makes our method fast. In the Supplementary, we show several examples of classification results, as a comparison. Figure 1(d) shows that with a bimodal online test-time adaptation (TTA) setup for CLIP, the proposed approach achieves a significant improvement compared to TPT and VTE.
Our main contributions are as follows:
@@
  • We conduct extensive experiments benchmarking the proposed method against established TTA baselines and others using CLIP on CIFAR-10C, CIFAR-100C, and ImageNet-C [18]. To generalize our method beyond common corruptions, we evaluate on various domain generalization datasets [14]-OfficeHome [51], PACS [25], VLCS [12], and Terra Incognita [3]. Overall, our approach achieves state-of-the-art performance for CLIP adaptation at test-time on various domain shifts.
@@ 3 Zero-shot performance analysis of CLIP to common image corruptions
where $\text{sim}(\cdot)$ is the cosine similarity and $\tau$ is the softmax temperature from CLIP’s pre-training stage. The text features $\mathcal{Z}$, in zero-shot evaluation, are always pre-computed. We also draw a comparison with CLIP performance on respective source test sets and follow this implementation  [^1] https://github.com/LAION-AI/CLIP_benchmark. Throughout this paper, the same TTA problem setup is used, unless mentioned otherwise. This TTA setup differs from the TTA setups of [36, 15, 34].
@@ 4 Proposed Methodology
The comprehensive analysis in Section 3 reveals that the zero-shot vision encoder $f_{vis}$ of CLIP is very sensitive to image corruption with increasing severity. Similarly, the performance of the text encoder $f_{txt}$ is invariant to the different text prompt templates and is also impractical to tune at test-time. Therefore, for effective adaptation, both the image and text encoders of CLIP need to be adapted to the incoming domain of test batches. We illustrate our proposed framework, BATCLIP, in Figure 4.
@@
Motivation of choice of parameters. Through thorough experimentation, Zhao et al. [58] demonstrate that fine-tuning LayerNorm (LN) parameters in attention-based models leads to yielding strong results on multimodal tasks. This also brings a reduction in trainable parameters. Updating normalization layers has also been deeply studied in tasks involving domain shifts [26]. Motivated by these works, we update the LayerNorm parameters of both CLIP encoders. This constitutes updating $\sim 0.044$% of all parameters. For every new task/corruption, we reset the model parameters following TENT [53] i.e., we perform a single domain TTA. In summary, our bimodal online test-time adaptation approach jointly updates the LayerNorm parameters of the CLIP encoders that are optimized synergistically through loss components aware of the input domain, leading to a more robust multimodal learning process.
@@ 5 Experiments and Results
Baselines. We compare our approach with zero-shot CLIP (ZS), state-of-the-art online TTA methods adopted for CLIP following [10] i.e., TENT [53], RoTTA [54], RPL [41], and SAR [35]. We update only the vision encoder $f_{vis}$ for such a setup. We benchmark against other online TTA methods using CLIP - TPT [45] and VTE [10]. Based on the discussion in Section 2, we adapt the two variants of WATT [36], for our setup, to WATT-P* and WATT-S*, while maintaining method consistency. For a “fairer” comparison, on each test batch, the model is adapted once using each of the suggested 8 prompt templates. Note that this setup is not online yet. In addition, CLIP parameters are reset after every task. We report the results on CIFAR-10C and CIFAR-100C only, based on their suggested settings. On ImageNet-C, we also report results against online StatA [55], based on their chosen hyperparameters. The model predictions can then be computed as in Eq. 1. We mention the details of each TTA method in the Supplementary.
Implementation Details. We query the text encoder $f_{txt}$ with a general prompt template “a photo of a $<$CLS$>$.” for all the datasets, as motivated earlier. We optimize both the vision encoder $f_{vis}$ and text encoder $f_{txt}$. For CIFAR-10C, we use an AdamW optimizer at a learning rate of $10^{-3}$. Similarly, for CIFAR-100C and ImageNet-C, Adam and AdamW optimizers are respectively used, at a fixed learning rate of 5$\times$$10^{-4}$, with the model being reset after each task. For fairness to all the baselines, the batch sizes are set to 200, 200, and 64 for the datasets, following various TTA benchmarks, at a corruption severity level of 5 for each task. For WATT-P* and WATT-S*, we use the same templates and learning rates as mentioned in [36]. For StatA, the class correlation in a batch is controlled by a Dirichlet distribution ($\gamma$). We report results for $\gamma$ = 0.1 (low correlation) and -1 (sequential). All experiments are run on a single NVIDIA RTX A5000 GPU using ViT-B/16 as the visual backbone. Results on ViT-B/32 are reported in the Supplementary.
@@ 5.2 Results: Domain Generalization Datasets
In this section, we evaluate our BATCLIP on commonly used domain generalization datasets [14] and benchmark against zero-shot ViT-B/16 (ZS), TENT, WATT-P* and WATT-S*. From Table 2, we observe that BATCLIP does better or comparable, across all datasets and domain shifts, to all the baselines. TENT matches zero-shot CLIP since LN normalization works per image across the features, independent of the batch statistics. We demonstrate that BATCLIP can be generalized and be robust to other various domain shifts, with a single step adaptation and global prompt template as opposed to WATT-P* and WATT-S* that use multiple prompt templates. The largest performance improvements come on the Terra Incognita dataset. The implementation details are described in the Supplementary.
@@ 5.3 Ablation Study and Analysis
Adaptation for multiple iterations. Here, we analyze adaptation for multiple iterations on a single batch, shown in Figure 6, using the ViT-B/16 backbone. Continuous adaptation of LN parameters, which normalize along the features of a single sample, can lead to overfitting as it fits to the noise patterns of the batch, ultimately degrading generalization to other domains. This could also lead to a decline in the loss of CLIP pre-trained knowledge. In the Supplementary, we report the post-adaptation results back on the source test sets - CIFAR10 and CIFAR100 to evaluate catastrophic forgetting. Our approach outperforms zero-shot CLIP, with a smaller decline in mean accuracy—especially on CIFAR-10C and ImageNet-C—showing that our loss components improve robustness under challenging conditions.
@@
$\mathcal{L}_{ent}$ + 0.1 * ($\mathcal{L}_{pm}$ + $\mathcal{L}_{sp}$) 71.90 41.88 30.24
$\mathcal{L}_{ent}$ + 0.5 * ($\mathcal{L}_{pm}$ + $\mathcal{L}_{sp}$) 73.70 41.98 30.44
@@
Ablation on loss components. In Table 3, we ablate the loss components by taking their different combinations using ViT-B/16. The task-wise results are reported in the Supplementary. We observe that the addition of $\mathcal{L}_{pm}$ and $\mathcal{L}_{sp}$ largely improve model performance than entropy minimization via $\mathcal{L}_{ent}$. In fact, the addition of $\mathcal{L}_{sp}$ brings larger improvements, proving that $f_{vis}$ produces discriminative features. Interestingly, using just $\mathcal{L}_{ent}$ achieves better or comparable accuracy than zero-shot CLIP, as the encoders work in synergy while adapting to the domain.
@@
Dataset TENT RPL RoTTA SAR Ours
ImageNet-C 40.62 40.00 39.76 42.10 42.84
(b) TTA baselines vs ours w/ ViT-L/14.
Table 4: Ablation on a larger backbone (ViT-L/14) - Mean accuracy (in %)
Extension to larger backbones. We run BATCLIP for a larger backbone like ViT-L/14 and report the results in Table 4. The key advantage lies in the projection matching loss ($\mathcal{L}_{pm}$) which updates the text encoder to produce more image-aware text features—an aspect missing in other state-of-the-art TTA baselines.
Computational efficiency. Due to an additional text encoder update, BATCLIP updates $\sim 0.044$% of all CLIP parameters, with a ViT-B/16 backbone, as opposed to $\sim 0.026$% in other TTA baselines. Though the increase is marginal, this leads to superior performances as adaptation of the text encoder produces input-aware CLIP text features, aligning better with the image features. On an NVIDIA RTX A5000 GPU, it takes 0.2s/batch on ImageNet-C with a batch size of 64, compared to 2.34s for WATT [36]. Also, WATT adapts per template (for several iterations), making it expensive every time. BATCLIP achieves faster on-the-fly adaptation rates and is favorable for real-time deployment.
@@ 6 Conclusion
    • In subsections 6.3.3 and 6.3.6, we present a detailed loss ablation study and the post-adaptation zero-shot generalization on source test sets, respectively — essentially evaluating catastrophic forgetting of BATCLIP.

### v3

@@ Abstract
Although open-vocabulary classification models like Contrastive Language Image Pretraining (CLIP) have demonstrated strong zero-shot learning capabilities, their robustness to common image corruptions remains poorly understood. Through extensive experiments, we show that zero-shot CLIP lacks robustness to common image corruptions during test-time, necessitating the adaptation of CLIP to unlabeled corrupted images using test-time adaptation (TTA). However, we found that existing TTA methods have severe limitations in adapting CLIP due to their unimodal nature. To address these limitations, we propose BATCLIP, a bimodal online TTA method designed to improve CLIP’s robustness to common image corruptions. The key insight of our approach is not only to adapt the visual encoders for improving image features but also to strengthen the alignment between image and text features by promoting a stronger association between the image class prototype, computed using pseudo-labels, and the corresponding text feature. We evaluate our approach on benchmark image corruption datasets and achieve state-of-the-art results in online TTA for CLIP. Furthermore, we evaluate our proposed TTA approach on various domain generalization datasets to demonstrate its generalization capabilities. Our code is available at https://github.com/sarthaxxxxx/BATCLIP.
@@ 1 Introduction
To address these core limitations, we propose, BATCLIP, a bimodal adaptation approach of CLIP for online TTA specifically, where both the visual and text encoders are jointly adapted by exploiting CLIP’s shared feature space of images and text. The overall objective is to achieve a strong alignment between image features and text features to enable more effective multimodal learning and adaptation. The adaptation procedure is two-fold: 1) Inspired by [55, 60] for efficient fine-tuning, we only adapt the norm layers, i.e., LayerNorm parameters of both encoders. However, such a model update does not consider the alignment of the encoder features. Therefore, to improve the alignment between class-specific visual and text features, we introduce a projection matching loss that maximizes the projection of the visual class prototypes with their corresponding text features. 2) To learn more discriminative visual features, we enhance the cosine distance between class prototypes, computed using pseudo-labels, to promote a more distinct separation in the image feature space. Our method is general-purpose, with no reliance on multiple prompt templates [37, 16, 35] or their ensembles, unlike VTE. We leverage batches of test samples for TTA, rather than focusing on single-image test adaptation, which makes our method fast. In the Supplementary, we show several examples of classification results, as a comparison. Figure 1(d) shows that with a bimodal online test-time adaptation (TTA) setup for CLIP, the proposed approach achieves a significant improvement compared to TPT and VTE. Our main contributions are as follows:
@@
  • We conduct extensive experiments benchmarking the proposed method against established TTA baselines and others using CLIP on CIFAR-10C, CIFAR-100C, and ImageNet-C [19]. To generalize our method beyond common corruptions, we evaluate on various domain generalization datasets [15]-OfficeHome [52], PACS [26], VLCS [13], and Terra Incognita [4]. Overall, our approach achieves SOTA performance for CLIP adaptation at test-time on various domain shifts (see Supplementary).
@@ 3 Zero-shot performance analysis of CLIP to common image corruptions
[…]
@@ 4 Proposed Methodology
The comprehensive analysis in Section 3 reveals that the zero-shot vision encoder $f_{vis}$ of CLIP is very sensitive to image corruption with increasing severity. Similarly, the performance of the text encoder $f_{txt}$ is invariant to the different text prompt templates and is also impractical to tune at test-time. Therefore, for effective adaptation, both the image and text encoders of CLIP need to be adapted to the incoming domain of test batches to focus on achieving strong image-text alignment and feature separability. We illustrate BATCLIP in Figure 4.
@@
Motivation of choice of parameters. Through thorough experimentation, Zhao et al. [60] demonstrate that fine-tuning LayerNorm (LN) parameters in attention-based models leads to yielding strong results on multimodal tasks. This also brings a reduction in trainable parameters. Updating normalization layers has also been deeply studied in tasks involving domain shifts [27]. Motivated by these works, we update the LayerNorm parameters of both CLIP encoders. This constitutes updating 0.044$\sim$ 0.044∼ 0.044% of all parameters. For every new task/corruption, we reset the model parameters following TENT [54] i.e., we perform a single domain TTA. In summary, our bimodal online test-time adaptation approach jointly updates the LayerNorm parameters of the CLIP encoders that are optimized synergistically through loss components aware of the input domain, leading to a more robust multimodal learning process.
@@ 5 Experiments and Results
Baselines. We compare our approach with zero-shot CLIP (ZS), state-of-the-art online TTA methods adopted for CLIP following [11] i.e., TENT [54], RoTTA [55], RPL [42], and SAR [36]. We update only the vision encoder $f_{vis}$ as in the baselines. We benchmark against other online TTA methods using CLIP - TPT [46] and VTE [11]. Based on the discussion in Section 2, we adapt the two variants of WATT [37], for our setup, to WATT-P* and WATT-S*, while maintaining method consistency. For a “fairer” comparison, on each test batch, the model is adapted once using each of the suggested 8 prompt templates. Note that this setup is not online yet. In addition, CLIP parameters are reset after every task. We report the results on CIFAR-10C and CIFAR-100C only, based on their suggested settings. On ImageNet-C, we also report results against online StatA [56], based on their chosen hyperparameters. The model predictions can then be computed as in Eq. 1. We mention the details of each TTA method in the Supplementary.
[…]
@@ 5.2 Results: Domain Generalization Datasets
In this section, we evaluate our BATCLIP on commonly used domain generalization datasets [15] and benchmark against zero-shot ViT-B/16 (ZS), TENT, WATT-P* and WATT-S*. From Table 2, we observe that BATCLIP does better or comparable, across all datasets and domain shifts, to all the baselines. TENT matches zero-shot CLIP since LN normalization works per image across the features, independent of the batch statistics. We demonstrate that BATCLIP can be generalized and be robust to other various domain shifts, with a single step adaptation and global prompt template as opposed to WATT-P* and WATT-S* that use multiple prompt templates. The largest performance improvements come on the Terra Incognita dataset. Large style variations in OfficeHome and PACS make it difficult for CLIP to generate effective text features, with a single step, from our default prompt. The implementation details are described in the Supplementary.
@@ 5.3 Ablation Study and Analysis
$\mathcal{L}_{ent}$ + $\mathcal{L}_{sp}$ 72.69 41.84 30.08
[…]
$\mathcal{L}_{ent}$ + $0.3(\mathcal{L}_{pm}$+$\mathcal{L}_{sp}$) 73.31 41.93 30.40
$\mathcal{L}_{ent}$ + $0.7(\mathcal{L}_{pm}$ +$\mathcal{L}_{sp}$) 73.85 42.04 30.61
@@
Adaptation for multiple iterations. Here, we analyze adaptation for multiple iterations on a single batch, shown in Figure 6, using the ViT-B/16 backbone. Continuous adaptation of LN parameters, which normalize along the features of a single sample, can lead to overfitting as it fits to the noise patterns of the batch, ultimately degrading generalization to other domains. This could also lead to a decline in the loss of CLIP pre-trained knowledge. In the Supplementary, we report the post-adaptation results back on the source test sets - CIFAR10 and CIFAR100 to evaluate catastrophic forgetting. Our approach outperforms zero-shot CLIP, with a smaller decline in mean accuracy, especially on CIFAR-10C and ImageNet-C, showing that our losses improve robustness. under challenging conditions.
Ablation on loss components. In Table 3, we ablate the loss components using ViT-B/16 (see more in the Supplementary). $\mathcal{L}_{pm}$ (row 4) and $\mathcal{L}_{sp}$ (row 5) improve model performance than entropy minimization via $\mathcal{L}_{ent}$. In fact, the addition of $\mathcal{L}_{sp}$ brings larger improvements, proving that $f_{vis}$ produces discriminative features. Interestingly, using only $\mathcal{L}_{ent}$ achieves comparable or better accuracy than zero-shot CLIP, with fully hyperparameter-free objectives.
@@
TTA Method ImageNet-C
TENT 40.62
RPL 40.00
RoTTA 39.76
SAR 42.10
Ours 42.84
(b) Results on ImageNet-C.
Table 4: Ablation on ViT-L/14 - Mean accuracy (in %).
Extension to larger backbones. We use ViT-L/14 and report the results in Table 4. The key advantage lies in the projection matching loss ($\mathcal{L}_{pm}$) which updates the text encoder to produce more image-aware text features—an aspect missing in other state-of-the-art TTA baselines.
Computational efficiency. Due to an additional text encoder update, BATCLIP updates 0.044$\sim$ 0.044∼ 0.044% of all CLIP parameters, with a ViT-B/16 backbone, as opposed to 0.026$\sim$ 0.026∼ 0.026% in other TTA baselines. Though the increase is marginal, this leads to superior performances as adaptation of the text encoder produces input-aware CLIP text features, aligning better with the image features. On an NVIDIA RTX A5000 GPU, it takes 0.2s/batch on ImageNet-C with a batch size of 64, compared to 2.34s for WATT [37]. Also, WATT adapts per template (for several iterations), making it expensive every time. BATCLIP achieves faster on-the-fly adaptation rates and is favorable for real-time deployment.
@@ 6 Conclusion
7 Acknowledgement
This project was supported by a grant from UT Dallas and an NVIDIA Academic Grant Program Award.
@@ 7 Acknowledgement
    • In subsections 7.3.3 and 7.3.8, we present a detailed loss ablation study and the post-adaptation zero-shot generalization on source test sets, respectively — essentially evaluating catastrophic forgetting of BATCLIP.
@@ 7.3.5 Impact of bimodal adaptation
Camera Sensor Condition ZS TENT SAR Ours
Auto-Exposure Light on 50.90 51.62 52.60 53.82
Light off 46.96 47.44 47.88 49.32
Manual Light on 60.44 60.64 60.30 61.26
Light off 60.76 60.99 59.62 61.61
Table 11: Online TTA experiments on ImageNet-ES [2]. Mean accuracy (in %).
SigLIP Clean (ImageNet) Source TENT SAR BATCLIP
ImageNet-C 82.00 35.44 37.58 39.62 40.10
Table 12: Online TTA experiments on ImageNet-C with SigLIP [58].
7.3.6 BATCLIP for other vision-language models
We employ SigLIP [58], a recent pre-trained vision-language model with 877 million parameters, for our online TTA experiments on ImageNet-C. We use a batch size of 8 and a learning rate of $5\times 10^{-5}$ with the AdamW optimizer. Despite the small batch size, which could impact prototypes, BATCLIP achieves better performance, as shown in Table 12.
7.3.7 Results on distribution shifts caused by lighting conditions, camera types, or object scales
To evaluate BATCLIP across a wide spectrum of shifts, we extend our setup to datasets exhibiting variations due to lighting conditions, camera types, and object scales. We conduct experiments on ImageNet-ES [2], which introduces significant variations in lighting and camera sensor settings (e.g., ISO, shutter speed). While more details can be found in the original work, but ImageNet-ES introduces wide variations in lighting conditions and camera sensor factors (ISO, shutter speed, etc.). We report the results in Table 11 using a ViT-B/16 backbone. To be noted that, “Auto-Exposure” has 5 tasks for each condition, while “Manual” has 27. On average, BATCLIP outperforms all the reported baselines.